OpenStudy (anonymous):

Define A = {a | a = n^2 for some n ∈ Z} and B = {b | b = m^6 for some m ∈ Z}. Prove that B ⊆ A.

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

I get that for n^2 and m^6 since m is in a higher power it will have the same even values that n will eventually have, but overall less values

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

do you just have to show that a^3 = b n^3 = a ?

OpenStudy (anonymous):

I guess that makes sense

OpenStudy (anonymous):

why n^3=a ?

OpenStudy (anonymous):

maybe better, is to say that any b can be expressed in terms of condition for a. I mean b=m^6=(k^3)^2 where k^3 is some n∈ Z

OpenStudy (anonymous):

an I think that's it

OpenStudy (unklerhaukus):

*ah typo,

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

Thanks Myko

OpenStudy (sirm3d):

to prove that \[B\subset A\] show that every \(b\in B\) implies \(b\in A\) let \(b\in B\) \(b=m^6\) for some \(m\in Z\) \[b=m^6=(m^3)^2\] put \(n=m^3\) \(b=m^6=n^2\in A\)